# Close-in compact super-Earth systems emerging from resonant chains: slow destabilization by unseen remnants of formation

### v1

@@
Systems of multiple planets with sizes between Earth and Neptune orbiting with short orbital periods have been discovered in the hundreds by transit surveys and are now known to be one of the most common types of planetary systems in the galaxy (Fressin2013; Petigura2013). These Kepler-like systems are increasingly understood to be a cohesive population with a set of correlated defining statistics, describing the physical properties of the planets and their orbital architectures (Weiss2018; VanEylen2015; He2020). Parameters such as planet multiplicity, orbital spacing, and typical planet sizes are well-understood for planets with radii between and and periods d. Formation models that aim to explain the origin of these systems must therefore focus on reproducing these statistics, rather than matching individual systems.
@@
Briefly, the typical Kepler-like system can be summarized as follows: the planets are either super-Earths with radii between 1 and , or sub-Neptunes above , the latter of which have a large volatile envelope. At the time of formation these super-Earths and sub-Neptunes are a single population, with the radius difference emerging later as a consequence of core mass and insolation-dependent atmosphere loss (Wu2019; Rogers2021; Owen2024). In terms of orbital architecture, many systems have multiple planets that are closely spaced in period but not near mean-motion resonance, while other systems have just a single transiting planet (Winn2015). A small number of systems have many planets locked in pairwise mean motion resonance; these are the “resonant chains.” Producing this diversity of orbital architectures for small planet systems is a major challenge of planet formation theory.
@@
A scenario that incorporates near-universal dynamical instabilities of primordial resonant chains is thus far the most successful model to reproduce the observed planetary census (Pu2015; Izidoro2017). In this model, resonant chains indeed form due to migration within the protoplanetary disk (Cresswell2006; Cresswell2008) but are catastrophically disrupted after the dissipation of the disk. During the instability, planetary orbits intersect and planets collide with each other, the star, or are ejected from the system entirely. Eventually, the instability ends and leaves behind a more widely spaced and dynamically excited system. The period-ratio distribution, which was strongly peaked near MMR during the disk phase, becomes smooth and nearly uniform, a close match to the data (Izidoro2017; Izidoro2021a). The model also explains the observed eccentricity distributions, including the anomalously eccentric single-transiting systems (VanEylen2019), patterns of intra-system uniformity in planet mass and period ratio (Goldberg2022), and lower densities and smaller mutual inclinations of near-resonant planets (Leleu2024).
@@
A major limitation of formation models incorporating N-body integrations is that they are restricted by computational constraints to the early stages of system evolution. Many transiting planet systems have Gyr or even 10 Gyr ages, corresponding to orbits of the innermost planet, a daunting challenge for even the fastest single-threaded integrators. Although substantial evolution takes place within the time covered by these models, much dynamical evolution can occur later. It is therefore not clear whether formation models that extend only to can be compared directly to data (Emsenhuber2021).
@@
A new, extremely powerful dataset is becoming available that promises to bridge this gap. The next generation of transit surveys, in particular TESS (Ricker2014), have discovered 10–20 multiplanet systems around stars younger than Gyr, and including several younger than 100 Myr (Hu2025). While still a small minority of the known exoplanet population, these young and adolescent systems are a unique window into the formation environment of small planets, otherwise unobservable with current methods. In a population study, Dai2024 showed that in the youngest systems ( Myr) mean-motion resonances are ubiquitous, appearing in of observed multiplanet systems. This fraction drops with a timescale of Myr to about for mature systems. Similar results have been corroborated by studies that derive stellar ages from galactic kinematics. Hamer2024 demonstrated that systems with planet pairs near resonance, particular second-order resonance, have smaller galactic velocity dispersions indicative of lower ages.
@@
On its face, this result appears to be direct confirmation of the “breaking-the-chains” hypothesis, because that idea predicts that systems emerging from their natal disks should be in resonant configurations that later break. On closer look however, many details are unresolved. For one, it has frequently been suggested that an additional mechanism outside of the N-body dynamics is necessary to break enough chains to match the observed distribution. Reshuffling of the system during disk dispersal has been proposed as one candidate process (Liu2017; Hansen2024; Pan2025), but this likely occurs too early to explain the abundance of near-resonance for systems with ages Myr. In fact, it is not obvious what astrophysical process can occur over these timescales.
@@
In this work, we demonstrate that resonant chains can unravel on timescales that match the observed decline in near-resonant frequency purely through N-body dynamics. In our model, planets grow out in a narrow ring (Batygin2023a; Ogihara2024) and then through Type I migration form resonant chains that resemble observed resonant systems. However, in most of them, chaotic diffusion gradually drives the system away from a fixed point until the resonant protection is lost and the planets undergo a dynamical instability. Importantly, the rate of this diffusion is frequently very slow, with a long tail, reproducing the observed frequency of near-resonance. Previous work has demonstrated that systems that begin resonant but later went unstable can readily match the transiting planet population (Izidoro2017; Izidoro2021a; Goldberg2022). Our primary goal in this work is to begin to understand how, why, and when this instability happens.
@@ 2 Methods
All simulations used the N-body code rebound with the trace integrator (Lu2024). trace is a hybrid time-reversible integrator that switches from the classical Wisdom-Holman Hamiltonian splitting (Rein2015) to direct Bulirsch-Stoer (BS) integration during close encounters. In comparison to the widely-used MERCURY code (Chambers1999) (and the similar rebound implementation mercurius (Rein2019)), trace is uniquely well-suited to this problem for several reasons. First, the switching distance is dynamically computed at each timestep rather than at the beginning of the integration, as it is in mercurius. While technically breaking symplecticity, this choice guarantees that the switching criteria remains robust even as planets grow or migrate (Hernandez2023; Lu2024). Secondly, trace properly resolves close pericenter approaches for highly-eccentric orbits, which can arise during dynamical instabilities. Finally, during close encounters, trace uses a much faster BS integrator than the IAS15 integrator used by mercurius, resulting in a major speed-up for very violent systems dominated by close encounters without a meaningful loss of accuracy. These benefits allowed us to perform significantly longer integrations and uncover slower instabilities.
[…]
@@
, , , , , and differs by simulation (Tab. 1). A constant aspect ratio was set to . This parametrization represents a disk with a pressure maximum at 1 au, and a transition from a 3/2 power-law profile in the outer disk to a typically flatter profile in the inner disk. Such a profile is frequently found in simulations of disk evolution with magnetic winds that efficiently carve the inner disk (Suzuki2016; Kunitomo2020). The disk was dissipated exponentially on a timescale of 1 Myr, and then at the disk was rapidly removed on a timescale of yr to represent disk photoevaporation. At the disk and its corresponding forces were removed completely. Each system was then integrated to under purely Newtonian gravity. The evolution of a typical simulation is shown in Fig. 1.
Embryos felt type I migration forces corresponding to Lindblad and corotation torques, computed using the equations of Paardekooper2010; Paardekooper2011. These formulae incorporate a dependence on the local power law index of surface density and temperature and account for saturation of the corotation torque, for which we use a Shakura-Sunyaev turbulent . We also included eccentricity and inclination damping forces according to the prescription of Cresswell2006; Cresswell2008, which account for reduction in damping at high inclination and eccentricity. For disks with , a migration trap at 1 au appears for a specific range of planet masses. The trap can hold planets in the outer system until they grow large enough to saturate the corotation torque, and quickly migrate inwards. While not derived from a specific physical model, this simulation setup is intended to represent a disk with a broad gas pressure bump and dust trap that triggers planetesimal and embryo formation (Flock2017).
The disk is also believed to be truncated at short radii when magnetic forces from the star are more important than accretion ram pressure. These sharp decreases in surface density also act as planet traps where inward migration is stopped and strong outward torques are possible (Masset2006). While some previous work prescribed a falloff in surface density at the inner edge and let corotation torques provide outward migration (e.g., Izidoro2017), Ataiee2021 showed that the standard torque prescription based on local power laws fails in these regimes. In fact, they found the trap to be much stronger than predicted, so that one planet at the inner edge is capable of halting the migration a resonant chain of several planets. We use a slightly modified version of their recommended prescription, in which we multiply the local migration rate by a coefficient
where , , , , au, and . This leads to a migration trap very close to , which is in very close agreement with the inner edge of Kepler systems being clustered near 10 d (Mulders2018). Fig. 2 shows a migration map at of this disk with the three values of used in this work.
@@ 3 Results
Following the disk dissipation, the system evolution takes one of two distinct paths. In the quiescent pathway, the resonant chain remains effectively unchanged over the entirety of the simulation, with purely periodic changes in orbital elements from resonant and secular evolution, reaching with librating resonant angles. On the other hand, many systems experience a dynamical instability sometime between and where the resonances become broken and the system is rapidly chaotic. In most cases, planetary orbits cross and planets collide very soon after the resonant angles begin circulating. The remaining systems are non-resonant, more widely spaced, more eccentric and inclined, and the planets are fewer and more massive. The quiescent pathway is interpreted as producing the rarely-observed resonant chains such as TRAPPIST-1 and TOI-178 (Gillon2017; Leleu2021), whereas the unstable track is responsible for the bulk of the known population of small planets in compact systems, which are non-resonant. The ability of a single process to reproduce both populations is the essence of the “breaking-the-chains” hypothesis originally proposed by Izidoro2017.
@@ 3.1 Planet accretion, migration, and resonance capture
The initial phase of planet accretion proceeds very quickly. Embryos quickly excite each other to crossing orbits and grow to several Earth masses. Typically, only 10–20 embryos remain by . In models with a pressure bump, embryos that reach the minimum mass to saturate the corotation torque migrate inwards quickly, sometimes resonantly entraining smaller planets that remain interior to them. Migration occurs earlier in straight disks, with planet masses set by a balance of local growth rates and the timescale to exit the ring (Batygin2023a). Migration always terminates at the inner edge located at 0.1 au, where the powerful planet trap prevents the chain from falling further towards the star.
The relative rates of migration for adjacent planets depends on their masses and the local disk profile, but convergent migration is frequently achieved because the inner planet is trapped by the inner edge directly or indirectly. Under these conditions, adjacent planet pairs capture efficiently into mean motion resonances, gradually building a large resonant chain anchored at the inner edge. Dynamical instabilities are repeatedly triggered during the disk phase either because migration dynamically overpacks the system or because of resonant overstability (Goldreich2014). However, high gas densities in the inner disk allow the broken chain to quickly reform. By the end of the disk phase at , 90–100% of systems are clearly resonant chains. The most common resonances are 5:3, 4:3, and 3:2 for the straight simulations and 4:3, 3:2, and 2:1 for 20flat and 40steep. Perhaps the biggest difference between the simulations is that those with a pressure bump frequently contain adjacent pairs of planets that are non-resonant: approximately 10% of pairs in 20flat and 40steep are further than 5% from a first- or second-order mean motion resonance, compared to in the straight simulations.
@@ 3.2 Population statistics
Figures 3–7 summarize the statistics of the synthetic population at the end of the 500 Myr integrations, divided by whether the system experienced an instability after the end of the disk phase. In these figures, all planets within 100 d have been included.
Some features are common to all simulation sets. Stable systems have more planets of lower mass, compared to unstable systems, and the period ratios of adjacent planets are dominated by resonant ratios. Most of these systems are extremely flat because their inclinations were damped by the protoplanetary disk and first-order resonances do not involve inclination, although a few have inclinations of . Eccentricities have been moderated excited by resonances to , a quantity which is set by the planet masses and balance of resonant excitation during migration vs. eccentricity damping in the disk (Terquem2019).
Unstable systems have qualitatively different features. In the closely-packed small planet regime, systems dynamically relax by way of collisions and mergers, thus reducing the number of planets but increasing the mass of the remaining ones. Prior resonances are systematically broken by the powerful planet-planet interactions during orbit crossing. The only source of energy dissipation is through inelastic mergers, which are insufficient to reform mean motion resonances. Final period ratios are determined by dynamical stability and dependent on planet masses and multiplicity (Pu2015). High mass runs thus have more widely spaced systems, from a median period ratio of 1.5 in 10straight to 2.49 in 40straight. Instabilities also leave eccentricities and inclinations in an excited state correlated with planet mass, although inelastic collisions sometimes leave systems anomalously flat or circular (Esteves2020).
The fraction of systems that are stable at the end of the 500 Myr integrations differs dramatically between simulation suites, from a low of 32.5% in 10straight to a high of 92.5% in 40steep. Two patterns are clear from Tab. 2. First, for the same disk profile, simulations with higher have a higher instability fraction. This is expected because over the regime of interest, the final resonance that planets capture into during migration depends much more strongly on the local gas surface density than on the planet masses (Kajtazi2023). High mass planets therefore capture into the same resonances as low mass planets, but are much more closely packed dynamically (Petit2020a; Goldberg2022a). The second clear trend is that, for the same , systems that formed in disks with a migration trap have a higher instability fraction. For example, 57.5% of systems in 20flat went unstable, compared to 35% in 20straight. Likewise, 40steep had a 92.5% unstable fraction compared to 80% in 40straight. We explore the possible origins of this trend in Section 4.1.

### v2

@@
Systems of multiple planets with sizes between Earth and Neptune orbiting with short orbital periods have been discovered in the hundreds by transit surveys and are now known to be one of the most common types of planetary systems in the galaxy (Fressin et al., 2013; Petigura et al., 2013). These Kepler-like systems are increasingly understood to be a cohesive population with a set of correlated defining statistics, describing the physical properties of the planets and their orbital architectures (Weiss et al., 2018; Van Eylen and Albrecht, 2015; He et al., 2020). Parameters such as planet multiplicity, orbital spacing, and typical planet sizes are well-understood for planets with radii between and and periods d. Formation models that aim to explain the origin of these systems must therefore focus on reproducing these statistics, rather than matching individual systems.
@@
Briefly, the typical Kepler-like system can be summarized as follows: the planets are either super-Earths with radii between 1 and , or sub-Neptunes above , the latter of which have a large volatile envelope. At the time of formation these super-Earths and sub-Neptunes are a single population, with the radius difference emerging later as a consequence of core mass and insolation-dependent atmosphere loss (Wu, 2019; Rogers and Owen, 2021; Owen and Schlichting, 2024). In terms of orbital architecture, many systems have multiple planets that are closely spaced in period but not near mean-motion resonance, while other systems have just a single transiting planet (Winn and Fabrycky, 2015). A small number of systems have many planets locked in pairwise mean motion resonance; these are the “resonant chains.” Producing this diversity of orbital architectures for small planet systems is a major challenge of planet formation theory.
@@
A scenario that incorporates near-universal dynamical instabilities of primordial resonant chains is thus far the most successful model to reproduce the observed planetary census (Pu and Wu, 2015; Izidoro et al., 2017). In this model, resonant chains indeed form due to migration within the protoplanetary disk (Cresswell and Nelson, 2006, 2008) but are catastrophically disrupted after the dissipation of the disk. During the instability, planetary orbits intersect and planets collide with each other, the star, or are ejected from the system entirely. Eventually, the instability ends and leaves behind a more widely spaced and dynamically excited system. The period-ratio distribution, which was strongly peaked near MMR during the disk phase, becomes smooth and nearly uniform, a close match to the data (Izidoro et al., 2017, 2021). The model also explains the observed eccentricity distributions, including the anomalously eccentric single-transiting systems (Van Eylen et al., 2019), patterns of intra-system uniformity in planet mass and period ratio (Goldberg and Batygin, 2022; Lammers et al., 2023), and lower densities and smaller mutual inclinations of near-resonant planets (Leleu et al., 2024).
@@
A major limitation of formation models incorporating N-body integrations is that they are restricted by computational constraints to the early stages of system evolution. Many transiting planet systems have Gyr or even 10 Gyr ages, corresponding to orbits of the innermost planet, a daunting challenge for even the fastest single-threaded integrators. Although substantial evolution takes place within the time covered by these models, much dynamical evolution can occur later. It is therefore not clear whether formation models that extend only to can be compared directly to data (Emsenhuber et al., 2021a).
@@
A new, extremely powerful dataset is becoming available that promises to bridge this gap. The next generation of transit surveys, in particular TESS (Ricker et al., 2014), have discovered 10–20 multiplanet systems around stars younger than Gyr, and including several younger than 100 Myr (Hu et al., 2025). While still a small minority of the known exoplanet population, these young and adolescent systems are a unique window into the formation environment of small planets, otherwise unobservable with current methods. In a population study, Dai et al. (2024) showed that in the youngest systems ( Myr), near-resonant configurations are ubiquitous, appearing in of observed multiplanet systems. This fraction drops with a timescale of Myr to about for mature systems. Similar results have been corroborated by studies that derive stellar ages from galactic kinematics. Hamer and Schlaufman (2024) demonstrated that systems with planet pairs near resonance, particular second-order resonance, have smaller galactic velocity dispersions indicative of lower ages.
@@
On its face, this result appears to be direct confirmation of the “breaking-the-chains” hypothesis, because that idea predicts that systems emerging from their natal disks should be in resonant configurations that later break. On closer look however, many details are unresolved. For one, it has frequently been suggested that an additional mechanism outside of the N-body dynamics is necessary to break enough chains to match the observed distribution. Reshuffling of the system during disk dispersal has been proposed as one candidate process (Liu et al., 2017; Hansen et al., 2024; Pan et al., 2025), but this likely occurs too early to explain the abundance of near-resonance for systems with ages Myr. In fact, it is not obvious what astrophysical process can occur over these timescales.
@@
In this work, we demonstrate that resonant chains can unravel on timescales that match the observed decline in near-resonant frequency purely through N-body dynamics. In our model, planets grow out in a narrow ring (Batygin and Morbidelli, 2023; Ogihara et al., 2024) and then through Type I migration form resonant chains that resemble observed resonant systems. However, in most of them, chaotic diffusion gradually drives the system away from a fixed point until the resonant protection is lost and the planets undergo a dynamical instability. Importantly, the rate of this diffusion is frequently very slow, with a long tail, reproducing the observed frequency of near-resonance. Previous work has demonstrated that systems that begin resonant but later went unstable can readily match the transiting planet population (Izidoro et al., 2017, 2021; Goldberg and Batygin, 2022). Our primary goal in this work is to begin to understand how, why, and when this instability happens.
@@ 2 Methods
All simulations used the N-body code rebound with the trace integrator (Lu et al., 2024). trace is a hybrid time-reversible integrator that switches from the classical Wisdom-Holman Hamiltonian splitting (Rein and Tamayo, 2015) to direct Bulirsch-Stoer (BS) integration during close encounters. In comparison to the widely-used MERCURY code (Chambers, 1999) (and the similar rebound implementation mercurius (Rein et al., 2019a)), trace is uniquely well-suited to this problem for several reasons. First, the switching distance is dynamically computed at each timestep rather than at the beginning of the integration, as it is in mercurius. While technically breaking symplecticity, this choice guarantees that the switching criteria remains robust even as planets grow or migrate (Hernandez and Dehnen, 2023; Lu et al., 2024). Secondly, trace properly resolves close pericenter approaches for highly-eccentric orbits, which can arise during dynamical instabilities. Finally, during close encounters, trace uses a much faster BS integrator than the IAS15 integrator used by mercurius, resulting in a major speed-up for very violent systems dominated by close encounters without a meaningful loss of accuracy. These benefits allowed us to perform significantly longer integrations and uncover slower instabilities.
The initial conditions of the simulations are summarized in Table 1. Following Batygin and Morbidelli (2023), who argue that planetary growth should commence from rings in which solids have concentrated, we initialize simulations at time by placing embryos of mass with semi-major axes drawn from a Gaussian distribution with mean 1 au and standard deviation around a central star of mass and radius . Their eccentricities and inclinations were drawn from a Rayleigh distribution with mode and , respectively, and their longitudes of pericenter, longitudes of ascending node, and mean longitudes were drawn from a uniform distribution from . All particles were set to be fully active. The timestep was set at the beginning of the integration to be the period of the innermost embryo, and periodically adjusted if it decreased below the period of a body due to inward migration. All bodies other than the star were set to have a physical radius corresponding to a density of and, if they approached closer than the sum of their radii, underwent perfect mergers conserving total mass and linear momentum.
@@
, , , , , and differs by simulation (Table 1). A constant aspect ratio was set to . This parametrization represents a disk with a pressure maximum at 1 au, and a transition from a 3/2 power-law profile in the outer disk to a typically flatter profile in the inner disk. Such a profile is frequently found in simulations of disk evolution with magnetic winds that efficiently carve the inner disk (Suzuki et al., 2016; Kunitomo et al., 2020). The disk was dissipated exponentially on a timescale of 1 Myr, and then at the disk was rapidly removed on a timescale of yr to represent disk photoevaporation. At the disk and its corresponding forces were removed completely. Each system was then integrated to under purely Newtonian gravity. The evolution of a typical simulation is shown in Fig. 1.
Embryos felt type I migration forces corresponding to Lindblad and corotation torques, computed using the equations of Paardekooper et al. (2010, 2011). These formulae incorporate a dependence on the local power law index of surface density and temperature and account for saturation of the corotation torque, for which we use a Shakura-Sunyaev turbulent . We also included eccentricity and inclination damping forces according to the prescription of Cresswell and Nelson (2006, 2008), which account for reduction in damping at high inclination and eccentricity. For disks with , a migration trap at 1 au appears for a specific range of planet masses. The trap can hold planets in the outer system until they grow large enough to saturate the corotation torque, and quickly migrate inwards. While not derived from a specific physical model, this simulation setup is intended to represent a disk with a broad gas pressure bump and dust trap that triggers planetesimal and embryo formation (Flock et al., 2017).
The disk is also believed to be truncated at short radii when magnetic forces from the star are more important than accretion ram pressure. These sharp decreases in surface density also act as planet traps where inward migration is stopped and strong outward torques are possible (Masset et al., 2006). While some previous work prescribed a falloff in surface density at the inner edge and let corotation torques provide outward migration (e.g., Izidoro et al., 2017), Ataiee and Kley (2021) showed that the standard torque prescription based on local power laws fails in these regimes. In fact, they found the trap to be much stronger than predicted, so that one planet at the inner edge is capable of halting the migration a resonant chain of several planets. We use a slightly modified version of their recommended prescription, in which we multiply the local migration rate by a coefficient
where , , , , au, and . This leads to a migration trap very close to 0.1 au, which is in very close agreement with the inner edge of Kepler systems being clustered near 10 d (Mulders et al., 2018). Figure 2 shows a migration map at of this disk with the three values of used in this work.
@@ 3 Results
Following the disk dissipation, the system evolution takes one of two distinct paths. In the quiescent pathway, the resonant chain remains effectively unchanged over the entirety of the simulation, with purely periodic changes in orbital elements from resonant and secular evolution, reaching with librating resonant angles. On the other hand, many systems experience a dynamical instability sometime between and where the resonances become broken and the system is rapidly chaotic. In most cases, planetary orbits cross and planets collide very soon after the resonant angles begin circulating. The remaining systems are non-resonant, more widely spaced, more eccentric and inclined, and the planets are fewer and more massive. The quiescent pathway is interpreted as producing the rarely-observed resonant chains such as TRAPPIST-1 and TOI-178 (Gillon et al., 2017; Leleu et al., 2021a), whereas the unstable track is responsible for the bulk of the known population of small planets in compact systems, which are non-resonant. The ability of a single process to reproduce both populations is the essence of the “breaking-the-chains” hypothesis originally proposed by Izidoro et al. (2017).
@@ 3.1 Planet accretion, migration, and resonance capture
The initial phase of planet accretion proceeds very quickly. Embryos quickly excite each other to crossing orbits and grow to several Earth masses. Typically, only 10–20 embryos remain by . In models with a pressure bump, embryos that reach the minimum mass to saturate the corotation torque migrate inwards quickly, sometimes resonantly entraining smaller planets that remain interior to them. Migration occurs earlier in straight disks, with planet masses set by a balance of local growth rates and the timescale to exit the ring (Batygin and Morbidelli, 2023). Migration always terminates at the inner edge located at 0.1 au, where the powerful planet trap prevents the chain from falling further towards the star.
The relative rates of migration for adjacent planets depends on their masses and the local disk profile, but convergent migration is frequently achieved because the inner planet is trapped by the inner edge directly or indirectly. Under these conditions, adjacent planet pairs capture efficiently into mean motion resonances, gradually building a large resonant chain anchored at the inner edge. Dynamical instabilities are repeatedly triggered during the disk phase either because migration dynamically overpacks the system or because of resonant overstability (Goldreich and Schlichting, 2014). However, high gas densities in the inner disk allow the broken chain to quickly reform. By the end of the disk phase at , 90–100% of systems are clearly resonant chains. The most common resonances are 5:4, 4:3, and 3:2 for the straight simulations and 4:3, 3:2, and 2:1 for 20flat and 40steep. Perhaps the biggest difference between the simulations is that those with a pressure bump frequently contain adjacent pairs of planets that are non-resonant: approximately 10% of pairs in 20flat and 40steep are further than 5% from a first- or second-order mean motion resonance, compared to in the straight simulations. Additionally, the coupling between mass and migration speed leads to a more pronounced mass ordering in the straight simulations so that the innermost planets are almost always the most massive. In contrast, in simulations with a migration trap, all planets interior to 0.5 au tend to have a similar mass.
@@ 3.2 Population statistics
Figure 3 summarizes the statistics of the 20flat synthetic population at the end of the 500 Myr integrations, divided by whether the system experienced an instability after the end of the disk phase. Similar figures for the other simulation suites are presented in Appendix A. In these figures, all planets within 100 d have been included.
Some features are common to all simulation sets. Stable systems have more planets of lower mass, compared to unstable systems, and the period ratios of adjacent planets are dominated by resonant ratios. Most of these systems are extremely flat because their inclinations were damped by the protoplanetary disk and first-order resonances do not involve inclination, although a few have inclinations of . Eccentricities have been moderately excited by resonances to , a quantity which is set by the planet masses and balance of resonant excitation during migration vs. eccentricity damping in the disk (Terquem and Papaloizou, 2019).
Unstable systems have qualitatively different features. In the closely-packed small planet regime, systems dynamically relax by way of collisions and mergers, thus reducing the number of planets but increasing the mass of the remaining ones. Prior resonances are systematically broken by the powerful planet-planet interactions during orbit crossing. The only source of energy dissipation is through inelastic mergers, which are insufficient to reform mean motion resonances. Final period ratios are determined by dynamical stability and dependent on planet masses and multiplicity (Pu and Wu, 2015). High mass runs thus have more widely spaced systems, from a median period ratio of 1.5 in 10straight to 2.49 in 40straight. Instabilities also leave eccentricities and inclinations in an excited state correlated with planet mass, although inelastic collisions sometimes leave systems anomalously flat or circular (Esteves et al., 2020).
The fraction of systems that went unstable before the end of the 500 Myr integrations differs dramatically between simulation suites, from a low of 32.5% in 10straight to a high of 92.5% in 40steep. Two patterns are clear from Table 2. First, for the same disk profile, simulations with higher have a higher instability fraction. This is expected because over the regime of interest, the final resonance that planets capture into during migration depends much more strongly on the local gas surface density than on the planet masses (Kajtazi et al., 2023). High mass planets therefore capture into the same resonances as low mass planets, but are much more closely packed dynamically (Petit et al., 2020; Goldberg et al., 2022). The second clear trend is that, for the same , systems that formed in disks with a migration trap have a higher instability fraction. For example, 57.5% of systems in 20flat went unstable, compared to 35% in 20straight. Likewise, 40steep had a 92.5% unstable fraction compared to 80% in 40straight. We explore the possible origins of this trend in Section 4.1.